OpenStudy (anonymous):

whats the square root of 15 X the square root of 12??

OpenStudy (anonymous):

Square root of 15 is 3.9 and the square root of 12 is 3.5.

OpenStudy (anonymous):

\[6\sqrt{5}\]

OpenStudy (cwrw238):

or sqrt(12*15) = sqrt 180

OpenStudy (cwrw238):

= sqrt 5*36 = 6 sqrt5

OpenStudy (anonymous):

Or 13.416407865

OpenStudy (cwrw238):

6 sqrt5 is the exact solution